OpenStudy (anonymous):

find the discriminant: x2 - 5x = 4 A) 3 B) 9 C) 18 D) 41

OpenStudy (accessdenied):

If we first put the equation into the form, \(ax^2 + bx + c = 0\), we can find the discriminant by using the formula and the coefficients: \(\text{discriminant} = b^2 - 4ac \). So first, can you put the equation in that form?

OpenStudy (accessdenied):

You recently sent me the message to subtract 4 from both sides. And I agree! x^2 - 5x - 4 = 0. Now we use the coefficients to compute this value: b^2 - 4ac. This is called the discriminant.